you must show all your work for full credit

Use the following information to answer the following questions

deductible: $1000

coinsurance: 20%

out-of-pocket maximum: $8000

let's say you have $30000 in covered medical expenses in January:

How much are you going to pay?

How much insurance company pays?

if you have an additional $20000 in medical expenses:

how much are you going to pay?

how much insurance company pays?
Answer #1

How much are you going to pay?

First, I need to pay amount equal to deductible i.e D

D=$1000

Amount left after deductibles=30000-1000=$29000

I am required to pay 20% of this amount i.e.CI

CI=20%*29000=$5800

Total payment required by me=D+CI=1000+5800=$6800

How much insurance company pays?

Amount to be paid by insurer=30000-6800=$23200

if you have an additional $20000 in medical expenses:

how much are you going to pay?

I am required to pay 20% of additional medical expenses, i.e CIa

CIa=20000*20%=$4000

I would be required to pay $4000 in addition.

how much insurance company pays?

Additional amount paid by insurance company=20000-4000=$16000